OpenStudy (anonymous):

A snail moved forward 4/7 inches during one hour, and 1/8 inches during the next hour. How much distance did it cover in the two hours? 39/56 in 25/56 in 5/56 in 3/56 in

OpenStudy (anonymous):

4/7 + 1/8 = (32+7)/(56) = 39/56 is the answer for this.

OpenStudy (anonymous):

4/7 + 1/8 = 39/56. Therefore, he moved that far.